OpenStudy (anonymous):

surface area of a cylinder with a base of 4yd and the height of 12yd

OpenStudy (lgbasallote):

is this total surface area or lateral surface area?

OpenStudy (anonymous):

the total surface area

OpenStudy (lgbasallote):

total surface area = 2(base area) + (base area)(height)...i am assuming 4 yd is the base area so it should be 4 yd^2...so TSA = 2(4) + (4)(12) =8 + 48 =56 yd^2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i'm not completely sure, but don't you need to take the perimeter of the base to multiply with the height? i'm sorry if i'm wrong and confusing you